OpenStudy (trojanpoem):

Find the limit:

OpenStudy (trojanpoem):

The limit of (d) of the complex number \[Z = \frac{ 1 }{ sind } + \frac{ 1 }{ cosd } i \] d -> pi/2 [No l'hopital]

OpenStudy (trojanpoem):

Are you thinking guys or the question doesn't show ?

OpenStudy (turingtest):

I was thinking, but now it seems you can just test the left and right hand limits and you sill see that the limit diverges

OpenStudy (trojanpoem):

Clarify a bit

OpenStudy (turingtest):

\[\large \lim_{d\to\pi/2^+}\frac1{\sin x}+\frac1{\cos x}i\\\large\lim_{d\to\pi/2^-}\frac1{\sin x}+\frac1{\cos x}i\]

OpenStudy (trojanpoem):

What is x ? and I am still confused.

OpenStudy (turingtest):

a typo, sorry\[\large \lim_{d\to\pi/2^+}\frac1{\sin d}+\frac1{\cos d}i\\\large\lim_{d\to\pi/2^-}\frac1{\sin d}+\frac1{\cos d}i\]

OpenStudy (turingtest):

do you know how to take a right-handed limit?\[\large \lim_{d\to\pi/2^+}\frac1{\sin d}+\frac1{\cos d}i\]

OpenStudy (trojanpoem):

Generally , I can. But in this one especially I can't.

OpenStudy (turingtest):

what is the limit of sinx as x approaches pi/2 from the positive side?

OpenStudy (trojanpoem):

sin(pi/2) = 1

OpenStudy (trojanpoem):

sinx/x if x approaches 0 Ops !

OpenStudy (turingtest):

sorry?

OpenStudy (trojanpoem):

Hmmm , seems like I am lost , aren't I ?

OpenStudy (turingtest):

perhaps lol do you know what a one-sided limit is?

OpenStudy (trojanpoem):

Yeah, It's when a function approaches a point from right or left ( higher and lower values).

OpenStudy (trojanpoem):

If that function has different left and right sided limits then It's limit doesn't exist.

OpenStudy (trojanpoem):

Am I right or still lost ?

OpenStudy (turingtest):

you are right, sorry i was distracted

OpenStudy (turingtest):

so as x approaches pi/2 from the right, what value does sinx have? what about cosx?

OpenStudy (trojanpoem):

You got the derivative ? " L'hopital rule" right ?

OpenStudy (turingtest):

you don't want to use l'hospital, right?

OpenStudy (trojanpoem):

Yep, written up !

OpenStudy (trojanpoem):

I don't want "L'hospital rule" as It's higher than my grade , and each time I ask for limit they use it.

OpenStudy (turingtest):

here is a unit circle

OpenStudy (turingtest):

as d approaches pi/2 from the right...is sind positive or negative?